# Learning to Synthesize Fault-Tolerant Quantum Circuits by Diagonalization

### v1

@@ 2 Qubits Heisenberg HHL Hubbard QAOA QPE Shor TFIM VQE Mean
Table II: Synthesis of two- and three- qubit unitaries taken from partitioned circuits. Compared to the QSD (top rows), the diagonalizing agent (bottom rows) on average reduces the number of $R_{Z}$ gates by 82.8% and 94.8% for two- and three- qubit unitaries respectively. Comparisons are made to the QSD because other synthesis tools fail to find solutions given $\epsilon=10^{-6}$.
Given a 2-3 qubit target, we wish to determine how well the DVNU based Reinforcement Learning diagonalization approach can synthesize unitaries. Throughout these experiments, we require that unitaries be synthesized to a distance of $\epsilon=10^{-6}$ (see Equation 2). At this level of precision, pure search-based tools are unable to synthesize unitaries by inversion. To test this assertion, we attempted to synthesize our suite of unitaries with Synthetiq: a simulated annealing tool which empirically achieves state-of-the-art performance [20]. We relaxed the requirement that Synthetiq fully invert matrices by using the diagonal distance function (Eq 6). Because Synthetiq attempts to invert matrices (i.e. uses a state representation in the form of $s_{t}=C_{t}U_{tar}^{\dagger}$ as opposed to $s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$), many targets remain out of reach. Solutions could only be found within a 20 minute time out period when inputs were already diagonal.
As a point of reference, we compare synthesis results of the diagonalizing agent to the Quantum Shannon Decomposition (QSD) [14]. This analytical decomposition was chosen because no other search-based agent that we tested was able find solutions by matrix inversion with sufficient precision. Results from the QSD are optimized by converting and combining $R_{Z}$ gates into Clifford+T gates when possible.
@@
We evaluated the diagonalizing agent’s performance on a set of unitaries taken from partitioned quantum algorithms. This suite of algorithms includes Shor’s Algorithm [44], TFIM, Heisenberg, and Hubbard model quantum chemistry simulation circuits [45], trained VQE [46] and QAOA [47] circuits, and Quantum Phase Estimation circuits [48]. The VQE and QAOA circuits were generated by MQTBench [49]. Each set of unitaries was filtered to ensure that every unitary was unique. There were 22,323 different 2-qubit unitaries and 45,202 different 3-qubit unitaries. Table II summarizes.
@@
We find that in the two-qubit case, diagonalization can find circuits implementing 42.9% of unitaries across all benchmarks. About 79% of these unitaries are already diagonal, therefore any synthesis agent using a diagonal distance (Equation 6) should find solutions in these cases. Most other unitaries contain 2-8 gates, typically Hadamard gates. These unitaries are often just controlled rotations about a non-Z axis.
@@
For the three-qubit case, realistic unitaries are far more complex: they are more likely to contain unitaries which do not conform to the diagonalization ansatz. Approximately 27.9% of three-qubit unitaries from our suite of partitioned circuits could be synthesized by the diagonalizing agent. Of these, approximately 57% were already diagonal. On average, 3.84 CNOT gates are needed to implement the diagonal operator, meaning the diagonalizing DVNU agent typically finds diagonalizing $L_{t}(\cdot)R_{t}$ implementations with an average of about 11 gates. Some common unitaries are almost trivial to synthesize under the diagonalizing approach. Toffoli (which is synthesized with optimal ancilla-free T-count and illustrated in Figure 3) and $CCR(\theta)$ gates are examples of such unitaries.
@@
Compared to the QSD decomposition, the diagonalizing agent produces solutions with far fewer non-trivial rotation gates. Implementing an $R_{Z}$ gate requires $O(\log\frac{1}{\epsilon})$ T gates when synthesized optimally with gridsynth [16]. For $\epsilon=10^{-6}$ this is approximately 60 T gates per $R_{Z}$. Lone T gates are almost negligible compared to continuous rotations in the high precision regime. Compared to the QSD, the DVNU synthesizer reduces the average number of non-Clifford gates by $82.8\%$ for two qubit unitaries and $94.8\%$ for three qubit unitaries.
@@
Although the success rate for the DVNU agent is lower than for mathematical decompositions, the potential savings in non-Clifford gates and the speed with the process runs remain strong arguments for its utility.
VI Discussion
Section V demonstrates how our model-based circuit synthesis agent performs. In this section, we discuss the implications of these results and pose future research directions.
@@ VI Discussion
(c) Proximity to the identity ($1-d_{HS}(s_{t},I)$) vs. gate count of random Clifford+T circuits. This value function is noisy and flat, making agents act randomly in most of the state space.
@@
Figure 9: The impact of the value function on state transitions. (a) Agents which use noisy and flat value functions to plan actions will act randomly. Nodes $A$ and $B$ represent the initial and terminal states respectively. Darker contours represent increasing value magnitudes. (b) Agents which use learned values to plan actions tend towards terminal states. (c) shows the Hilbert-Schmidt distance which is a noisy and flat value function. Lines are the mean distances of 250 random circuits, the shaded regions are interquartile ranges. In synthesis with discrete gate sets, unitary distance does not correlates with gate count. States which require fewer gates to reach a terminal state generally do not correspond to lower distances than states which require more gates. (d) shows the learned value function which is based on path distance. States which require fewer gates to reach a terminal state have higher values than those which require more gates. Gate cancellation in randomly generated circuits contributes to variance in state values.
VI-A Trade-Offs for High Precision Synthesis of Realistic Unitaries
Unitary synthesis has proven to be an invaluable tool for transpiling near-term quantum circuits [50]. Synthesis algorithms targeting Fault Tolerant gate sets are certain to fit a similar niche. In order to be useful for transpilation, FT synthesis algorithms must be able to approximate unitaries to high levels of precision. In this paper, we have demonstrated how our Reinforcement Learning based diagonalization approach can synthesize realistic unitaries to a precision of $\epsilon=10^{-6}$. This value was chosen as it is 1000×1000$\times$1000 × more precise than previous methods (see Table I). Because the state representation is much more complex in the diagonalization approach ($s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$ instead of $s_{t}=C_{t}U_{tar}^{\dagger}$ ), the search process is much more difficult. To this end, the modifications made to the neural architecture of DVNU (Section IV) are essential for machine learning approaches to perform well. Further innovations may further improve the applicability of deep learning to datasets of unitary matrices.
Because much of the difficulty, and the need for approximation, comes from translating $R_{Z}$ rotations to Clifford+T gates, our diagonalizing tool can produce implementations of many unitaries even more precise than this. Targeting higher levels of precision using DVNU often requires retraining, which takes approximately 2-3 days to complete if done from scratch.
Assuming that there are continuous rotations that must be very precisely approximated may not always be necessary. In some cases, as in early demonstrations of FT hardware for example, lower levels of precision (on the order of $\epsilon=10^{-3}$) may be acceptable. In these scenarios, pure search-based tools such as Synthetiq are a better choice. Exhaustively searching the space of circuit implementations may be viable in this case, and it is likely to produce lower gate count results.
Determining exactly what precision is needed to ensure circuits produce meaningful results is problem dependent, and an area that deserves deeper investigation. Analytical decompositions can always be used to produce higher precision approximations, but because these decompositions require so many gates, new techniques must be invented to handle this task. We expect that other analytical-search-based hybrid approaches are possible and practical.
VI-B Interpreting the Value Function
@@
In order for a function to meaningfully guide search-based synthesis, it should correlate with the goal of finding short gate count implementations of unitaries. Intuitively, the value of a unitary state should depend on the number of gates needed to transform that state to a terminal state.
@@
A discrete gate set synthesis algorithm which relies directly on the Hilbert-Schmidt distance function to guide the placement of gates will randomly search the space of possible circuits. Figure 9(c) shows how for random Clifford+T circuits, Hilbert-Schmidt distance to the identity does not correlate with the number of gates in the circuit (and therefore the number of gates needed to synthesize that unitary). Synthesis states in the neighborhood of the current state appear equally distant from a terminal state.
@@
How does a Reinforcement Learning agent avoid this pitfall? Under a reward function such as Equation 7, the learned value of states correlates with the number of gates in the circuit. As shown in Figure 9(d), the value of a state is a learned, scaled and, shifted representation of the “gates to go”. This property makes the agent a fast and efficient synthesizer; it ideally takes the shortest direct path to a terminal state.
@@
In the case of exact synthesis of single-qubit Clifford+T unitaries, the smallest divisor exponent (SDE) plays a similar role as the value function [12]. This synthesis algorithm sequentially takes actions depending on which gate decreases the SDE. In the multi-qubit case, it is believed that such a simple function does not naturally exist [17]. The model-based RL approach taken in this paper can be construed as a learned version of the SDE function for approximate multi-qubit synthesis.
VI-C Scaling to Wider and Deeper Circuits
Only cases of 1-3 qubit synthesis are considered in this paper. Circuits of interest are typically on the order of 100-1000s of logical qubits. As the size the unitaries associated with these circuits grows exponentially in the number of qubits, it is not likely that circuits larger than 10s of qubits wide will ever be directly synthesizable. Past works have addressed this issue with circuit partitioning; by splitting large circuits into smaller subcircuits which can each be individually synthesized [51, 52, 53, 54]. Circuit partitioning is applicable when a circuit implementation of a program is already known. This is a fair assumption to make in the case of gate set transpilation and optimization. Our experiments in Section V have demonstrated that an RL diagonalizing agent is capable of synthesizing high precision unitaries taken directly from partitioned circuits of interest.
In non-discrete synthesis scenarios (i.e. with a get set of parameterized single-qubit unitaries and CNOTs as opposed to Clifford+T), numerical methods have proved invaluable to finding compact circuit implementations of unitaries [55, 56, 57]. Recent work has investigated how these methods can be transferred to the setting of Clifford+T synthesis [58]. As this approach requires solving a complex non-linear constrained optimization problem, its run time can quickly grow unwieldy; synthesis with DVNU requires on the order of seconds as opposed to tens of minutes. As our diagonalization method runs much faster, we foresee that it may be used as a “first pass” optimization method in conjunction with these more powerful synthesizers; cases that cannot be easily diagonalized can instead be handled by constrained numerical optimization or analytical decompositions.
Our method requires only a representation of the state to determine values. This means that the value of placing composite gates can be determined as easily as gates directly represented in the gate set. Scaling to deeper circuits may therefore be possible if a set of common composite gates are collected.
VII Conclusion

### v2

@@ 2 Qubits Heisenberg HHL Hubbard QAOA QPE Shor TFIM VQE Mean
Non-Clifford Reduction 95.8% 93.2% 97.6% 92.7% 96.8% 90.4% 94.5% 98.3% 95.1%
Table 2. Synthesis of 2- and 3-qubit unitaries taken from partitioned circuits (see Figure 10 for an illustration). Numbers indicate average time, success rate, Clifford, and non-Clifford gate counts for unitaries taken from a variety of benchmarks. Compared to the QSD (top rows), the diagonalizing agent (bottom rows) on average reduces the number of $R_{Z}$ gates by 83.5% and 95.1% for 2- and 3-qubit unitaries respectively. Comparisons are made to the QSD because other synthesis tools fail to find solutions given $\epsilon\leq 10^{-4}$.
Given a 2-3 qubit target, we wish to determine how well the DVNU-based Reinforcement Learning diagonalization approach can synthesize unitaries. Throughout these experiments, we require that unitaries be synthesized to a distance of $\epsilon=10^{-6}$ (see Equation 2). In many cases, the diagonalizing agent produces far more precise circuit implementations of unitaries. In Section 5.4, we synthesize unitaries to an even higher precision. At this value of $\epsilon$ pure search-based tools are unable to synthesize unitaries by inversion. To test this assertion, we attempted to synthesize our suite of unitaries with Synthetiq: a simulated annealing tool which empirically achieves state-of-the-art performance (Paradis et al., 2024). Because Synthetiq attempts to invert matrices, many targets remain out of reach. Solutions could only be found within a 20 minute time out period when inputs contained no continuous rotations (i.e. $U_{3}$ or $R_{Z}$ gates).
As a point of reference, we compare synthesis results of the diagonalizing agent to the Quantum Shannon Decomposition (QSD) (Shende et al., 2006). This analytical decomposition was chosen because no other search-based tool that we tested was able find solutions with sufficient precision. Results from the QSD are optimized by converting and combining $R_{Z}$ gates into Clifford+T gates when possible.
@@
We evaluated the diagonalizing agent’s performance on a set of unitaries taken from partitioned quantum algorithms (see Figure 10 for an illustration of how circuits are partitioned). This suite of algorithms includes Shor’s Algorithm (Shor, 1997), TFIM, Heisenberg, and Hubbard model quantum chemistry simulation circuits (Bassman et al., 2022), trained VQE (Peruzzo et al., 2014) and QAOA (Farhi et al., 2014) circuits, and Quantum Phase Estimation circuits (Kitaev, 1995). The VQE and QAOA circuits were generated by MQTBench (Quetschlich et al., 2023). Each set of unitaries was filtered to ensure that every unitary was unique. There were 22,323 different 2-qubit unitaries and 45,202 different 3-qubit unitaries. Table 2 summarizes.
Unitaries which are diagonal, or nearly diagonal, are ubiquitous primitives in realistic quantum benchmarks. These common unitaries are nearly trivial to synthesize under the diagonalizing approach, but often are out of reach for inversion-based synthesis algorithms. The Toffoli gate (which is synthesized with optimal ancilla-free T-count and illustrated in Figure 4) and $CCR(\theta)$ gates are examples of commonplace, nearly diagonal, unitaries.
@@
We find that in the two-qubit case, diagonalization can find circuits implementing 42.9% of unitaries across all benchmarks. About 79% of these unitaries are already diagonal, and therefore trivial for a diagonalizing agent to synthesize. Most other unitaries contain 2-8 gates, typically single-qubit Hadamard and S gates.
@@
For the three-qubit case, realistic unitaries are far more complex: they are more likely to contain unitaries which do not conform to the diagonalization ansatz (see Figure 4). Approximately 27.9% of three-qubit unitaries from our suite of partitioned circuits could be synthesized by the diagonalizing agent. Of these, approximately 57% were already diagonal. On average, 3.84 CNOT gates are needed to implement the diagonal operator, meaning the diagonalizing DVNU agent typically finds diagonalizing $L_{t}(\cdot)R_{t}$ implementations with an average of about 11 Clifford+T gates.
@@
Compared to the QSD, the diagonalizing agent produces solutions with far fewer non-trivial rotation gates. Implementing an $R_{Z}$ gate requires $O(\log\frac{1}{\epsilon})$ T gates when synthesized optimally with gridsynth (Ross and Selinger, 2016). For $\epsilon=10^{-6}$ this is approximately 60 T gates per $R_{Z}$. Lone T gates are almost negligible compared to continuous rotations in the high precision regime. Compared to the QSD, the DVNU synthesizer reduces the average number of non-Clifford gates by $83.5\%$ for two qubit unitaries and $95.1\%$ for three qubit unitaries.
@@
Although the success rate for the DVNU agent is lower than for mathematical decompositions, the potential savings in non-Clifford gates and the speed with the process runs remain strong arguments for its utility. Next, we demonstrate how the DVNU approach to synthesis can be used to transpile future-term quantum algorithms into the Clifford+T gate set.
5.4. A Fault-Tolerant Circuit Synthesis Workflow
Unitary synthesis algorithms are well suited for transpiling quantum circuits into new sets of gates. Past work  (Younis and Iancu, 2022) targeting NISQ gates has demonstrated how multi-qubit synthesis techniques can produce implementations of higher quality than when using optimal single- and two-qubit (Tucci, 2005) synthesis. In this scenario, the additional optimization power comes from replacing gate-by-gate local optimization with more global optimization of multi-qubit unitaries.
Our DVNU agent is able to find solutions to a far more general set of unitaries at high precision than the current state-of-the-art methods. The question remains whether in the FT case techniques that are able to handle complex unitaries are better than composing optimal techniques for simple (lower dimension) unitaries.
To demonstrate how diagonalization can be leveraged to transpile algorithms into FT gate sets with synthesis we compare the results of multi-qubit FT synthesis to that of gate-level transpilation; where circuits are transpiled gate-by-gate (as opposed to subcircuit-by-subcircuit) into the Clifford+T gate set. This strategy is used as a point of comparison for the DVNU agent because the only other methods that produce high precision implementations of these unitaries, the QSD, results in $O(4^{n})$ T gates (Table 2).
Figure 10. Fault-Tolerant gate set transpilation using unitary synthesis. Quantum algorithms are partitioned into many subcircuits. These subcircuits are transpiled independently using unitary synthesis. The optimized and transpiled subcircuits are then replaced into the original circuit. $R_{Z}$ gates are handled in a post-processing stage using gridsynth.
By Gate Add HHL Mult QAE QFT QPE Shor
Qubits 17 6 16 50 32 30 16
$T$ Gates 22,646 21,977 97,250 60,750 97,231 28,890 73,446
Cliffords 35,404 34,243 152,188 97,048 152,375 45,867 117,492
$\epsilon_{\text{total}}$ $0.3\times 10^{-6}$ $0.2\times 10^{-6}$ $1.1\times 10^{-6}$ $0.7\times 10^{-6}$ $1.1\times 10^{-6}$ $0.3\times 10^{-6}$ $0.8\times 10^{-6}$
2Q Blocks $T$ Gates 21,656 21,797 97,250 58,014 97,232 27,987 74,446
Cliffords 33,867 34,194 152,188 92,675 152,268 44,348 117,492
$\epsilon_{\text{total}}$ $0.6\times 10^{-6}$ $0.3\times 10^{-6}$ $1.2\times 10^{-6}$ $4.8\times 10^{-6}$ $4.4\times 10^{-6}$ $1.1\times 10^{-6}$ $0.8\times 10^{-6}$
% Diagonalized 94.0% 72.4% 92.7% 45.1% 91.0% 21.4% 26.2%
Improvement $4.4\%$ $0.8\%$ $0\%$ 4.5% $0\%$ 3.1% $0\%$
3Q Blocks $T$ Gates 19,489 21,707 80,935 59,760 85,441 28,440 73,446
Cliffords 30,748 33,857 127,048 96,999 135,051 45,361 117,492
$\epsilon_{\text{total}}$ $0.4\times 10^{-6}$ $0.3\times 10^{-6}$ $1.2\times 10^{-6}$ $4.2\times 10^{-6}$ $3.1\times 10^{-6}$ $1.0\times 10^{-6}$ $0.8\times 10^{-6}$
% Diagonalized 85.1% 14.3% 73.5% 40.6% 85.3% 14.4% 12.5%
Improvement 13.9% 1.2% 16.8% $1.6\%$ 12.1% $1.6\%$ $0\%$
Table 3. Transpilation results. The By Gate strategy indicates gate-level transpilation into Clifford+T gates. The 2Q and 3Q Block strategies indicate circuits partitioned into subcircuits of that size and synthesized by diagonalization. In scenarios where diagonalization fails, subcircuits are transpiled gate by gate. Each subcircuit is synthesized to a distance of $\epsilon=10^{-8}$. Individual $R_{Z}$ gates are synthesized to $\epsilon=10^{-9}$, resulting in at least 90 $T$ gates per $R_{Z}$. The total approximation error in each circuit on the order of $\epsilon_{\text{total}}\approx 10^{-6}$.
A high level summary of the process used is shown in Figure 10. For 3-qubit synthesis, our method is as follows:
  1. (1)
    Partition a quantum circuit into 3-qubit blocks that contain as many gates as possible.
  2. (2)
    For each 3-qubit unitary, use the DVNU agent to generate a high precision approximation using diagonalization.
  3. (3)
    If DVNU fails to synthesize a unitary, split it into 2-qubit unitaries and again attempt synthesis.
  4. (4)
    For all unitaries that still do not meet the threshold, take the original partitioned subcircuit and convert single-qubit parameterized gates into a sequence of $R_{Z}$ and Clifford gates. Use gate replacement rules for other non-Clifford+T gates.
  5. (5)
    Decompose $R_{Z}$ gates with gridsynth produced Clifford+T sequences.
Table 3 summarizes these results across several quantum algorithms and primitives. The total approximation error $\epsilon$ across the entire circuit is on the order of $10^{-6}$. This value is an upper bound that is found by summing the individual approximation errors for each partition and each gridsynth transpiled $R_{Z}$ gate (Wang and Zhang, 1994). Transpilation via diagonalization results in savings in the number of T gates needed for these algorithms compared to the gate-level transpilation strategy. We see the highest T gate count reduction (16.8%) for the 16 qubit multiplier circuit. Other algorithmic primitives such as the 17 qubit adder circuit, and the 32 qubit QFT see similarly large reductions. The HHL, QAE, and QPE circuits see more moderate decreases in T count.
The performance of the diagonalizing agent for transpilation is highly dependent on the circuit partitioning algorithm, which in this case is driven by circuit structure. For example, a 16-qubit implementation of Shor’s algorithm sees no improvement when using diagonalization compared to the gate-level transpilation strategy. The 2-qubit diagonalizing agent is only able to find transpiled implementations to 26% of unitaries taken from this circuit. We likewise observe that only 12% of partitioned 3-qubit subcircuits are diagonalized for this same circuit. Shor’s algorithm consists of many repeated copies QFT and inverse QFT modules, which we know can be simplified well with diagonalization. For 2- and 3-qubit partitions, 91% and 85% of subcircuits taken from the example QFT circuit can be transpiled by the DVNU agent. This indicates that performance of FT transpilation algorithms which leverage synthesis are highly dependent on the partitioning algorithm used.
Combined with circuit partitioning, unitary synthesis enables transpilation to take place on a less localized scale. This more global view of a circuit’s function allows for synthesis to outperform simple gate-level rule-based methods.
6. Discussion
Resource efficient transpilation into FT gate sets is vital for practical quantum error corrected computation. To this end, high precision synthesis targeting the Clifford+T gate set is a invaluable asset. Here, we discuss how diagonalization-based synthesis with DVNU can be further developed in pursuit of this goal.
6.1. Trade-Offs for High Precision Synthesis of Realistic Unitaries
In order to be useful for large scale FT computation, synthesis algorithms must be able to approximate unitaries to high levels of precision. In the worst case, the total approximation error in a transpiled circuit is the sum of each individual gate’s or partition’s approximation error (Wang and Zhang, 1994). If the circuits transpiled in Section 5.4 (see Table 3) had instead been handled by a synthesis algorithm capable of only finding solutions to a precision of $\epsilon=10^{-3}$, the average total approximation error of the transpiled circuits would be upper bounded by $\epsilon_{\text{total}}=0.56$ (ranging from 0.05 to the maximum error of 1.0). This amount of approximation error is unlikely to lead to meaningful algorithm outputs. The outlook for coarse synthesis is even worse for larger width circuits.
Because diagonalization allows for $12^{n}-$ 1 continuous rotations to be handled analytically, much higher levels of precision are possible with this approach. In fact, we can consider the class of exactly diagonalizable unitaries (analogous to exactly synthesizable unitaries), where $U=RD_{\theta}^{-1}L$ and each entry of $L,R\in\mathbb{Z}[e^{i\pi/4},\frac{1}{2}]$. These exactly diagonalizable unitaries can be approximated to arbitrarily high levels of precision using synthesis by diagonalization, so long as implementations can be found for $L$ and $R$. Common examples of these unitaries include those that can be constructed with sequences of controlled rotation gates $CR(\theta)$.
We have demonstrated how our Reinforcement-Learning-based diagonalization approach can synthesize realistic unitaries to precisions between $\epsilon=10^{-6}$-$10^{-8}$. These values are orders of magnitude ($10^{3}$-$10^{5}\times$) more precise than previous methods (see Table 1) and enable high precision transpilation of complete algorithms (see Table 3). Determining the exact precision needed to ensure circuits produce meaningful results is an open area of research in both NISQ and FT domains.
No matter the end-precision required, this work showcases the benefit of augmenting analytical decomposition methods (which result in many non-Clifford gates) with search-based multi-qubit methods. Although diagonalization produces higher quality results than pure analytical decompositions, these results are likely non-optimal in non-Clifford gate counts. We expect that other analytical-search-based hybrid approaches that offer further improvements are both possible and practical.
@@ 6.1. Trade-Offs for High Precision Synthesis of Realistic Unitaries
(c) Proximity to the identity ($1-d_{HS}(s_{t},I)$) vs. gate count of random Clifford+T circuits. This value function is noisy and flat, causing agents to act randomly in most of the state space.
@@
Figure 11. The impact of the value function on state transitions. (a) Agents which use noisy and flat value functions to plan actions will act randomly. Nodes $A$ and $B$ represent the initial and terminal states respectively. Darker contours represent increasing value magnitudes. (b) Agents which use learned values to plan actions tend towards terminal states. (c) shows the Hilbert-Schmidt distance which is a noisy and flat value function. Lines are the mean distances of 250 random circuits, the shaded regions are interquartile ranges. In synthesis with discrete gate sets, unitary distance does not correlates with gate count. States which require fewer gates to reach a terminal state generally do not correspond to lower distances than states which require more gates. (d) shows the learned value function which is based on path distance. States which require fewer gates to reach a terminal state have higher values than those which require more gates. Gate cancellation in randomly generated circuits contributes to variance in state values.
6.2. Interpreting the Value Function
@@
In order for a function to meaningfully guide search-based synthesis, it should correlate with the goal of finding short gate count implementations of unitaries. Intuitively, the value of a synthesis state should depend on the number of gates needed to transform that state into a terminal state.
@@
A discrete gate set synthesis algorithm which relies directly on the Hilbert-Schmidt distance function (or Frobenius norm, diamond norm, etc.) to guide the placement of gates will randomly search the space of possible circuits. Many synthesis algorithms to date use these cost functions to guide gate placement (Davis et al., 2020; Smith et al., 2023; Paradis et al., 2024). Figure 11(c) shows how for random Clifford+T circuits, Hilbert-Schmidt distance does not correlate strongly with the number of gates in the circuit (and therefore the number of gates needed to synthesize that unitary). Synthesis states in the neighborhood of the current state appear equally distant from a terminal state.
@@
How does a Reinforcement Learning agent avoid this pitfall? Under a reward function such as Equation 7, the learned value of states correlates with the number of gates in the circuit. As shown in Figure 11(d), the value of a state is a learned, scaled and, shifted representation of the gates to go. This property makes the agent a fast and efficient synthesizer; it ideally takes the shortest direct path to a terminal state.
@@
In the case of exact synthesis of single-qubit Clifford+T unitaries, the smallest divisor exponent (SDE) plays a similar role as the value function (Kliuchnikov et al., 2012). This synthesis algorithm sequentially takes actions depending on which gate decreases the SDE. In the multi-qubit case, it is believed that such a simple function does not naturally exist (Gosset et al., 2013). The model-based RL approach taken in this paper can be construed as a learned version of the SDE function for approximate multi-qubit synthesis.
6.3. The Utility of Diagonalization for Compilation Tasks
By augmenting search-based compilation with generalized analytical decomposition, our DVNU agent is able to greatly expand the domain of unitaries that can be transpiled and increases the precision to which they are transpiled. However, diagonalization also enables even more powerful compilation strategies.
The use of ancilla qubits and projective measurements exposes more opportunity for optimization when implementing operations in FT gate sets. Probabilistic Quantum Circuits with Fallback (Bocharov et al., 2015a) and Repeat Until Success (Bocharov et al., 2015b) schemes leverage these resources to reduce the number of non-Clifford gates needed to implement $R_{Z}$ rotations compared to optimal ancilla-free synthesis. These techniques rely on synthesizing unitaries with a particular structure (the so called Jack of Daggers structure). As a powerful synthesis approach, diagonalization can be used to explore other structures which can systematically decrease non-Clifford gate counts when paired with ancilla.
By loosening the synthesis objective from inversion to diagonalization, our tool also exposes opportunity for resource-efficient implementations of continuous rotations. (Cody Jones et al., 2012) leverages phase kickback and adder circuits to express single-qubit $Z$-axis rotations much more efficiently than a standard gate decomposition. Our technique naturally finds multi-qubit $Z$-axis rotations (diagonal unitaries). It may be possible to combine these techniques to more efficiently approximate complex continuous rotations across multiple qubits.
As most FT synthesis work has focused on the Clifford+T gate set, there may be unexplored optimization opportunity when considering alternate gate sets such as Clifford+$\sqrt{T}$ and Clifford+V. Search-based synthesis tools are well suited to begin answering these questions, as synthesis in alternative gate sets is simply a matter of modifying the set of actions an agent can apply.
7. Conclusion
We have demonstrated a novel approach to high precision unitary synthesis targeting fault-tolerant gate sets. By diagonalizing unitaries instead of directly inverting them, complex continuous degrees of freedom can be avoided then handled efficiently using mathematical decomposition methods. This enables us to synthesize a broader scope of unitaries than other search-based synthesis methods which directly invert continuous rotations, a task which is intractable for high levels of precision. The diagonalization process is general and other synthesis tools can be retrofitted to diagonalize rather than invert unitaries.